OpenStudy (anonymous):

help factoring trinomials. x2 + 14x + 49

mathslover (mathslover):

\[\large{x^2+14x+49}\] what r the factors of 49 ? @ahoymewmew

mathslover (mathslover):

@ahoymewmew please reply

OpenStudy (anonymous):

You might like to find A and B where: A*B = 1*49=49 A+B = 14